OpenStudy (anonymous):

Use the three steps to solve the problem. Two years ago a father was four times as old as his son. Three years from now the father will be only three times as old as the son. How old is each now? {son is yrs., father is yrs.}

OpenStudy (ajprincess):

Let the present age of son be x and the age of father be y. Can u tell me what was the age of the father two yrs ago?

OpenStudy (anonymous):

no i dont know i am srry

OpenStudy (ajprincess):

Subtract 2 from the present age y.

OpenStudy (anonymous):

so it would be 2-y*4 right?

OpenStudy (ajprincess):

hmm no. The age of father two yrs ago is y-2. And the son,s is x-2. Since it is given that the father's age was 4 times the age of son y-2=4(x-2). Getting it?

OpenStudy (anonymous):

kinda

OpenStudy (ajprincess):

Is it confusing? If so plz ask.

OpenStudy (anonymous):

yes it's confusing

OpenStudy (ajprincess):

How old are u nw?

OpenStudy (ajprincess):

what was ur age before two yrs?

OpenStudy (ajprincess):

U subtracted to frm ur present age to find the age bfore 2 yrs. right?I assumed father's present age to be y. then his age before two yrs wud have been y-2. Is is it clear?

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

The age of the son bfore two yrs wud hav been x-2. Getting it?

OpenStudy (anonymous):

mmhhh

OpenStudy (ajprincess):

sorry? isn't t clear?

OpenStudy (anonymous):

it is

OpenStudy (ajprincess):

Since it is given that the age of the father was 4 times the son's age y-2=4(x-2)

OpenStudy (ajprincess):

Getting it?

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

What will be the age of the father after three years?

OpenStudy (anonymous):

y+3

OpenStudy (ajprincess):

yup right:)

OpenStudy (ajprincess):

What will be the age of the son?

OpenStudy (anonymous):

x+3

OpenStudy (ajprincess):

Since it is given that the father's age will be three times of the age of the son y+3=3(x+3) getting it?

OpenStudy (anonymous):

okay

OpenStudy (ajprincess):

So u hav two eqns y-2=4(x-2) y+3=3(x+3) Can u find x and y?

OpenStudy (anonymous):

y-2=4x-8 y+3=3x=9

OpenStudy (ajprincess):

y-2=4x-8 y+3=3x+9 ya right:). Nw solve these equations to find x and y. can u do it?

OpenStudy (anonymous):

idk

OpenStudy (ajprincess):

y-2=4x-8 what do u get when u add 2 to both sides?

OpenStudy (anonymous):

10

OpenStudy (ajprincess):

y-2+2=4x-8+2 y=4x-6

OpenStudy (ajprincess):

y+3=3x+9 What do u get when u subtract 3 from both of the sides?

OpenStudy (anonymous):

12

OpenStudy (ajprincess):

hmm no. y+3-3=3x+9-3 y=3x+6 Getting it?

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

So u have two eqns y=4x-6 y=3x+6. Since both 4x-6 and 3x+6 are equal to y you can equate them. 4x-6=3x+6 Can u find x nw?

OpenStudy (anonymous):

10

OpenStudy (ajprincess):

hmm no. 1.Add 6 both sides 2.Subtract 3x both sides. What do u get nw?

OpenStudy (anonymous):

1x

OpenStudy (ajprincess):

That is vat u get in left hand side. What do u get in ur right hand side?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

@robtobey can u help me?

OpenStudy (anonymous):

Solve the following for f and s, the father's and son's current age respectively.\[\left\{\frac{f-2}{s-2}=4,\frac{f+3}{s+3}\text{=}3\right\}\]Can you solve the above for f and s?

OpenStudy (anonymous):

idk i dont know how to do it

OpenStudy (anonymous):

Is the following easier?\[\{f-2=4(s-2),f+3=3(s+3)\}\]

OpenStudy (anonymous):

i dont know this is all confusing to me

OpenStudy (anonymous):

focus on the equation on the left. f - 2 = 4 (s - 2) Multiply out the RHS f - 2 = 4 s - 8 Move -2 to the RHS. f = 4 s - 8 + 2 Add -8 and +2 f = 4 s -6 Replace f in the second equation with 4 s -6 f + 3 = 3 (s + 3) (4 s -6) + 3 = 3 (s + 3) Can you solve the above for s and get a number?

OpenStudy (anonymous):

son is 7

OpenStudy (anonymous):

I have to log off now. The answer is\[\{f=42,s=12\} \] You should verify the answer yourself.

OpenStudy (anonymous):

okay